A human (88 kg) goes skydiving from space and goes into a vertical dive trying to break their speed record. How fast are they traveling if they fall for a full mile while in "space" aka no air resistance?
Answer #1

Given :

m = 88 kg

initial speed (u) = 0 m/s

distance traveled (d) = 1 mile = 1609.34 m

Using formula : v2 - u2 = 2ad

Here, a = g = 9.81 m/s2

v2 - u2 = 2gd

v2 - (0)2 = (2)(9.81 m/s2)(1609.34 m)

v2 = 31575.2508 m2/s2

v = 177.69 m/s